OpenStudy (anonymous):

(m - 5) − 16 = −12 m

TheSmartOne (thesmartone):

1. Add `6 to both sides

TheSmartOne (thesmartone):

16***

OpenStudy (uri):

LETS SEE WHO WINS

OpenStudy (uri):

#fail.

TheSmartOne (thesmartone):

@pootypie @~ Can you do that? lol

TheSmartOne (thesmartone):

There are so little questions that everyone is jumping on them.. lol

TheSmartOne (thesmartone):

And what are we doing solving for m? @pootypie

OpenStudy (uri):

@TheSmartOne wins @Nnesha loses

TheSmartOne (thesmartone):

``` (m - 5) − 16 = −12 m ``` I don't see any solve for m...

OpenStudy (igreen):

Okay..well first we add 16 to both sides..-12 + 16 = ?

OpenStudy (anonymous):

I Got 4? 16+ -12 = 4 And 5 + 16= 21

OpenStudy (igreen):

Yep, that gives us: m - 5 = 4 Add 5 to both sides, what's 5 + 4?

TheSmartOne (thesmartone):

Well if remove the parenthesis